Solve for C:
0.884 = (pi/2)[(c)^2 - (c/2)^2]

Answer by user_dude2008(1862)

0.884 = (pi/2)[(c)^2 - (c/2)^2]


0.884 = (pi/2)[c^2 - (c^2)/4]


0.884 = (pi/2)c^2 - (pi/2)(c^2)/4


8*0.884 = 4pi*c^2 - pi*c^2


7.07200 = 4pi*c^2 - pi*c^2


7.07200 = 3pi*c^2


c=+-sqrt(7.07200/3pi)